why does 3 minus 5 minus 6 multiply by 8 equals minus 50
Answer by ewatrrr(24785)

 
Hi,
why does 3 minus 5 minus 6 multiply by 8 equals minus 50
(3-5) - 6·8 = -2 - 48 = -50